OpenStudy (anonymous):

Suppose an automobile has a kinetic energy of 2600 J. When it moves with two times the speed, what will be its kinetic energy? Answer in units of J

ganeshie8 (ganeshie8):

\(KE = \frac{1}{2} mv^2\)

ganeshie8 (ganeshie8):

when v becomes 2v :- \(KE_{new} = \frac{1}{2} m(2v)^2\) \(~~~~~~ = 4 \times \frac{1}{2} mv^2\) \(~~~~~~ = 4 \times KE\)

OpenStudy (anonymous):

thank you :)

ganeshie8 (ganeshie8):

np :)